OpenStudy (anonymous):

ind the vertex, value of p, axis of symmetry, focus, and directrix of the parabola. y−2=1/12(x−4)^2

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

y-y1=(1/12)(x-x1)^2

OpenStudy (gorv):

vertex (x1,y1)

OpenStudy (gorv):

what is vertex ??

OpenStudy (gorv):

line of symmetry x=x1

OpenStudy (gorv):

at @fani1996

OpenStudy (anonymous):

hmmm (4,2)

OpenStudy (anonymous):

focus( 4,5)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (perl):

do you have general equation of a parabola with vertex

OpenStudy (anonymous):

idk im confused

OpenStudy (perl):

General form of parabola with vertex (h,k), there are two cases: (x-h)^2 = 4p (y - k ) or (y-k)^2 = 4p(x-h)

OpenStudy (perl):

in your problem, the square is on the (x-h) , so we will use the first form

OpenStudy (anonymous):

hmm ok

OpenStudy (anonymous):

the axis of symmerty x= -1?

OpenStudy (perl):

General form of parabola with vertex (h,k).There are two cases: Case 1: Parabola opens upwards or downwards equation: (x-h)^2 = 4p (y - k ) focus: (h,k+p) directrix : y = k-p Axis : x = h -------------------------------------------------------- Case 2: Parabola opens to the right or left (sideways) equation: (y-k)^2 = 4p(x-h) Focus: (h+p, k) Directrix: x = h- p Axis : y = k

OpenStudy (anonymous):

are my answers right

OpenStudy (anonymous):

do i have the focus and vertex

OpenStudy (perl):

Given problem : 1/12(x-4)^2 = y-2 Get into form above (x-4)^2 = 12 (y-2) now it is in form (x-h)^2 = 4p(y-k) h = 4, k = 2 , 4p = 12 ==> p = 3

OpenStudy (perl):

ok so far?

OpenStudy (anonymous):

hmm yes i had p=3

OpenStudy (perl):

focus: (4,2+3) directrix : y = 2-3 Axis : x = 3

OpenStudy (perl):

diretrix : y = -1

OpenStudy (anonymous):

axis of symmetry x= 4

OpenStudy (anonymous):

ahh so i had the vertex and focus and p= 3 right thanks for helping me and i understand better